find the values of x and y solve the fllowing system of equations:
-4x+3y=17
-7x+9y=11

Answer by amoati(11)

-4x+3y=17 -- (1)
-7x+9y=11 -- (2)
multiplying (1) by 3 gives -12x+9y=51 then this one minus(2) gives:
-5x=40 --> x = -8 putting this is in (1) gives y = -5